OpenStudy (anonymous):

Solve for r: C = 2pir. r = C over 2 pi r = 2 c all over pi r = C over pi r = pi times C all over 2

OpenStudy (shamil98):

\[\huge C = 2{\pi}r\] \[\huge \frac{ C }{ 2{\pi} } = r\]

OpenStudy (anonymous):

so the answer is C. Thanks~

OpenStudy (shamil98):

The answer is A, that is C/2pi not C/pi lol

OpenStudy (anonymous):

Kayy Thanks!